OpenStudy (anonymous):

solve sqrt6q-5=5 answer 25?

OpenStudy (anonymous):

no, Idek how you got that, 6q-5=25 6q=30 q=5

OpenStudy (anonymous):

me either bro lol

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

q= 50/3

OpenStudy (anonymous):

wait, is it the sqrt(6q-5) or sqrt(6q) -5 ???

OpenStudy (anonymous):

sqrt (6q-5) = 5

OpenStudy (anonymous):

yeah, Ireices ur wrong.